On Saturday evening, the slave laser switched off while. The system was in standby for the weekend ; this morning we went to EER to investigate the issue.
The safety chain had been triggered due to a faulty temperature controller, namely the controller of diode 1.
We installed a spare controller. As usual, the programming procedure is not straightforward; additional information for possible future interventions is provided at the end of this entry.
We were eventually able to restart the slave laser. We gradually increased the current of the pumping diodes and returned to a relatively normal operating condition. The output power of the Neovan is slightly lower than usual, and the mismatch in the PMC is slightly worse (Normalized reflected power 5% higher). We will perform fine tuning tomorrow.
In addition to the information already provided in entries 66004 and 62780, we add the following details:
-
The communication addresses are:
-
diode 1: 27
-
diode 2: 1
-
xtal 1: 2
-
xtal 2: 3
-
-
To access the third menu (PID parameter settings), one must press the right buttons simultaneously, then when “Ulock” appears one should use either 0, 1, or 10 (the value depends on user access rights).
-
At present, the controllers appear to be configured with the following Ulock values (tbc)
-
diode 1: 1
-
diode 2: 0
-
xtal 1: 10
-
xtal 2: 10
-
As observed previously, the safety chain reset did not initially work because the start button on EER1/C3 also needed to be pressed. It is unclear whether this action was possible from the beginning or only after off/on of the 6 V power supply (the unit with the red switch).